Integrative-Risk and valuation Giant Enterprises stock has a required return of 14.8%. The company, which plans to pay a divYear Dividend per Share $2.45 $2.28 $2.10 2019 2018 2017 $1.95 2016 2015 $1.82 $1.80 $1.73 2014 2013 Please show how you solved
Answer #1

a. The risk premium on the stock is = Required return - risk free rate

= 14.8 % - 4%

= 10.8%

The historical growth rate of dividends is,

1.73* (1 + g) ^6 = $2.45

So, the growth rate of dividends is

g = 5.9721%

g = 6%

b. The constant growth model. the value of stock is,

Po = D1/(Re - g)

= $2.6/ ( 0.148 - 0.06)

= $29.5455

= $29.55 ( rounded off )

A decrease in risk premium , would lower the required return and which would increase the price of the stock.

Re = Rf + beta ( Rm - Rf)

If the risk premium is reduced, the required rate of return will also reduce.

Po = D1/ Re - g

So, as the required return decreases, the stock price increases. As the stock price and the required return are inversely related.